# Generalizable Image Repair for Robust Visual Autonomous Racing

### v1

@@ Abstract
Vision-based autonomous racing relies on accurate perception for robust control. However, image distribution changes caused by sensor noise, adverse weather, and dynamic lighting can degrade perception, leading to suboptimal control decisions. Existing approaches, including domain adaptation and adversarial training, improve robustness but struggle to generalize to unseen corruptions while introducing computational overhead. To address this challenge, we propose a real-time image repair module that restores corrupted images before they are used by the controller. Our method leverages generative adversarial models, specifically CycleGAN and pix2pix, for image repair. CycleGAN enables unpaired image-to-image translation to adapt to novel corruptions, while pix2pix exploits paired image data when available to improve the quality. To ensure alignment with control performance, we introduce a control-focused loss function that prioritizes perceptual consistency in repaired images. We evaluated our method in a simulated autonomous racing environment with various visual corruptions. The results show that our approach significantly improves performance compared to baselines, mitigating distribution shift and enhancing controller reliability.
@@
Video: https://youtu.be/o8EpQvsltPE
@@ I INTRODUCTION
Vision-based autonomous driving has achieved significant progress in racing applications [1]. Despite these advancements, real-world deployment remains challenging due to image distribution shifts [2] caused by sensor noise, adverse weather conditions (such as rain, fog, and snow), rapid lighting changes, and adversarial perturbations. Such factors can lead to severe image corruption, compromising the stability and reliability of vehicle control systems. In practice, most existing vision-based methods assume that the training data accurately reflects the test environment. However, this assumption often fails, resulting in subpar decision-making and degraded driving performance.
@@
This paper addresses the critical challenge of maintaining control performance in the presence of corrupted images. We introduce a real-time image repair module that seamlessly integrates into vision-based driving systems. This module repairs degraded images at test time without any prior knowledge of them, thereby reducing the distribution gap between the training and testing environments. Our method leverages recent advances in GANs, specifically CycleGAN [9] and pix2pix [10], to perform real-time image restoration. CycleGAN enables unpaired image-to-image translation, allowing adaptation to unseen corruptions without requiring paired training data, while pix2pix utilizes paired data to learn a direct mapping between corrupted and clean images, leading to more precise restorations when such data is available.
By utilizing these techniques, our approach balances computational efficiency and repair quality, ensuring minimal latency while preserving essential perception features for reliable control. To train our module, we introduce a control-focused loss function that ensures the repaired images lead to accurate control outputs, directly improving autonomous racing performance under adverse conditions.
@@
  • A real-time image repair module that integrates into vision-based autonomous racing systems, enhancing robustness against image corruption.
@@ II PROBLEM STATEMENT
  • Vision-based control module $h:\mathbf{X}\rightarrow\mathbf{Y}$, which takes an image as input and is realized by a neural network
@@
Given a training dataset $\mathcal{D}_{train}$ of paired images and control actions, where the images are drawn i.i.d from the image distribution $P_{train}$ and the actions follow the center line of the track, we define an image-based neural network controller $h$, which translates image observations $y$ into control actions $u$. During operation, we have a testing image distribution $P_{test}$, where images may undergo corruptions and disturbances. To formalize the relationship between the testing and training images, we assume that test images in $P_{test}$ are generated from clean images in $P_{train}$ via an unknown random noise $\eta$ from an unknown distribution $P_{\eta}$. Then we consider two settings of image generation:
@@
We define an image repair model, $r:\mathbf{Y}\rightarrow\mathbf{Y}$, that translates images from the testing domain to the training domain without explicit knowledge of the corruption functions $c$ and $C$ or noises $\eta$ or $P_{\eta}$. Intuitively, the repair model aims to recover an image $y^{\prime}=r(\hat{y})$ to move the restored image closer to the training distribution $P_{train}$, thereby improving controller performance. Our architecture is shown in Figure 2.
@@
For the autonomous car racing setting, we calculate this difference in trajectory by comparing mean-squared cross-track error (CTE) values with respect to the center line of the track. Specifically, the corruptions we consider include darkness, salt-and-pepper noise, rain, fog, and snow, which we denote as $\eta\in\{\eta_{\text{dark}},\eta_{\text{salt}},\eta_{\text{rain}},\eta_{\text{fog}},\eta_{\text{snow}}\}$. Samples of these corruptions are shown in Figure 1.
@@ III-A Repair Techniques
Adversarial restoration: Much like in adversarial training, many restoration methods opt for learning-based approaches. For instance, PixelDefend [15] and generative adversarial denoising [16] reconstruct perturbed images to restore clean data representations, randomized smoothing [17] certifies robustness by averaging predictions over noisy variations of an input, and feature denoising networks [18] integrate adversarial noise suppression within convolutional architectures. However, these methods also come with increased computational complexity and their efficacy often depends on the strength and nature of adversarial attacks.
@@
where $r_{\text{LR}}$ is the Lucy Richardson Deconvolution approach, $r_{\text{VB}}$ is the Variational Bayesian approach, $y$ is the original image from $P_{test}$, and $y^{\prime}$ is the repaired image that is then passed to the controller as follows:
@@
We implement three generative approaches — VAEs, CycleGAN, and pix2pix — and compare their performance in Section IV. We prepend our generative image repair model $r$ to $h$, repairing corrupted images $\hat{y}\in P_{test}$ and passing them to $h$:
@@ III-B Generative Model Training
For pix2pix, the training parallels that of CycleGAN, but with the requirement having a paired dataset of normal and corrupted images. We train for 30 epochs and then take the generator as our repair model. Much like for CycleGAN, we implement an additional controller loss $\mathcal{L}_{GS}$, representing the L1 loss between the predicted control action on normal images $y$ and predicted images $y^{\prime}$ from the generator. The loss is then added to the existing pix2pix generator loss:
@@ IV EXPERIMENTAL EVALUATION
Config Setup Normal Brightness Salt/Pepper Rain Fog Snow Unseen All
@@
- Lucy Richardson Deconvolution 2.957 3.148 2.459 2.088 1.969 1.923 2.362 2.471
- Variational Bayes 2.158 3.078 2.552 2.733 2.523 2.275 2.646 2.571
@@
VAE 1.395 3.147 2.088 1.838 1.942 2.220 2.085 2.171
CycleGAN (Unpaired, tr: $\varnothing$, te: $h$) 1.611 1.499 2.559 1.424 1.178 1.965 1.620 1.764
CycleGAN (Unpaired, tr: $\varnothing$, te: $h_{adv}$) 1.282 1.525 1.726 1.533 1.475 2.306 1.936 1.673
CycleGAN (Paired, tr: $h$, te: $h$) 1.242 1.353 1.345 1.392 1.087 1.833 1.507 1.394
CycleGAN (Paired, tr: $h$, te: $h_{adv}$) 1.519 1.392 2.234 1.396 2.347 2.364 2.355 1.927
CycleGAN (Paired, tr: $h_{adv}$, te: $h$) 1.466 1.709 1.885 1.405 2.069 1.956 2.013 1.765
CycleGAN (Paired, tr: $h_{adv}$, te: $h_{adv}$) 1.502 1.763 1.602 1.663 1.480 2.320 1.946 1.745
pix2pix (Paired, tr: $\varnothing$, te: $h$) 1.648 1.403 1.552 1.451 1.415 2.003 1.734 1.605
pix2pix (Paired, tr: $\varnothing$, te: $h_{adv}$) 1.588 1.767 1.628 2.323 1.464 2.154 1.842 1.847
pix2pix (Paired, tr: $h$, te: $h$) 1.984 1.950 1.966 1.922 1.602 2.028 1.936 1.951
pix2pix (Paired, tr: $h$, te: $h_{adv}$) 1.635 1.552 2.261 2.312 2.361 2.276 2.319 2.095
pix2pix (Paired, tr: $h_{adv}$, te: $h$) 1.817 1.846 1.931 1.997 2.121 1.959 2.042 1.948
pix2pix (Paired, tr: $h_{adv}$, te: $h_{adv}$) 1.655 2.189 1.534 2.149 2.254 2.124 2.190 2.004
@@
VAE 1.393 3.147 2.146 1.923 1.935 2.071 1.998 2.167
CycleGAN (Unpaired, tr: $\varnothing$, te: $h$) 1.500 1.446 1.723 1.264 1.266 2.189 1.788 1.597
CycleGAN (Unpaired, tr: $\varnothing$, te: $h_{adv}$) 1.405 1.419 1.516 1.639 2.239 2.106 1.887 1.752
CycleGAN (Paired, tr: $h$, te: $h$) 1.570 1.506 1.117 1.513 1.338 1.995 1.770 1.530
CycleGAN (Paired, tr: $h$, te: $h_{adv}$) 1.507 1.567 1.412 1.668 1.639 2.075 1.883 1.658
CycleGAN (Paired, tr: $h_{adv}$, te: $h$) 1.265 2.332 2.782 2.580 2.830 1.869 2.253 2.343
CycleGAN (Paired, tr: $h_{adv}$, te: $h_{adv}$) 1.455 2.628 2.265 2.297 2.702 2.166 2.233 2.289
pix2pix (Paired, tr: $\varnothing$, te: $h$) 2.652 2.250 1.853 1.586 1.673 1.865 1.731 2.013
pix2pix (Paired, tr: $\varnothing$, te: $h_{adv}$) 1.594 1.719 1.676 1.888 2.124 2.071 1.981 1.856
pix2pix (Paired, tr: $h$, te: $h$) 1.840 1.548 1.540 1.529 1.454 1.919 1.784 1.656
pix2pix (Paired, tr: $h$, te: $h_{adv}$) 2.063 1.614 1.643 1.809 1.761 2.085 1.952 1.838
pix2pix (Paired, tr: $h_{adv}$, te: $h$) 1.796 1.804 1.682 1.574 1.406 1.341 1.523 1.643
pix2pix (Paired, tr: $h_{adv}$, te: $h_{adv}$) 1.801 1.875 1.743 1.817 1.621 2.066 1.946 1.826
Experimental Setup. The primary objective is to assess how well different repair techniques restore corrupted images and improve control performance under various disturbances. To evaluate the effectiveness of our image repair methods in vision-based autonomous racing, we conduct experiments in a simulated environment for DonkeyCar racing (see https://www.donkeycar.com/). Simulation is chosen over real-world deployment due to its controlled setting, reproducibility, ability to systematically introduce diverse corruptions, and an easy-to-quantify autonomous racing task.
On the other hand, open datasets like nuScenes and Waymo Open do not allow counterfactual executions under visual corruptions, whereas implementing repeatable paired corruptions in a physical setting can be very effortful. The implementation of our evaluation can be found in an online repository: https://github.com/Trustworthy-Engineered-Autonomy-Lab/generalizable-image-repair.
@@
Our proposed methods include CycleGAN and pix2pix. For CycleGAN, we experiment with both unpaired and paired training with a controller loss. The unpaired CycleGAN learns a mapping between $P_{test}$ (corrupted images) and $P_{train}$ (clean images) without explicit supervision. In contrast, the paired CycleGAN is trained with corrupted-clean image pairs and incorporates an additional loss function to align repaired images with optimal control actions. Finally, we train a pix2pix model on paired corrupted-clean images, both with and without a controller loss, to examine its effectiveness compared to CycleGAN-based methods.
@@
Overall, these results confirm the effectiveness of integrating generative repair models for mitigating visual corruptions, as illustrated in our video online. Notably, incorporating a controller loss during CycleGAN training significantly enhances repair quality, leading to more stable and reliable control under challenging conditions. This highlights the importance of aligning image repair objectives with downstream control performance, rather than relying solely on visual reconstruction quality.
@@ V-C Adversarial Robustness
For robustness against adversarial perturbations, adversarial training (AT) [36] uses adversarial examples in training. Methods like FGSM [5], PGD [11], and TRADES [12] minimize discrepancies between clean and perturbed predictions. AT has also been extended to reinforcement learning [13] and domain adaptation [14] to improve robustness. However, as previously discussed, AT still struggles to generalize to adversarial examples outside of the training set. While the controller performs well on the three known corruptions, it struggles to generalize to the other two unseen noises.
Instead of modifying training, adversarial purification removes perturbations from input images. PixelDefend [15] and generative adversarial denoising [16] reconstruct images, while randomized smoothing [17] improves robustness by averaging predictions over perturbed inputs. Feature denoising networks [18] integrate noise suppression directly in convolutional architectures, reducing adversarial vulnerabilities. To compare our method to purification approaches, we tested Lucy Richardson Deconvolution and Variational Bayes, which proved detrimental to controller performance on both normal and corrupted images. Thus, classical computer vision approaches struggle to handle complex and unstructured noises, such as those seen in the real world (rain, fog, snow).

### v2

# Generalizable Image Repair for Robust Visual Control

@@ Abstract
Vision-based control relies on accurate perception to achieve robustness. However, image distribution changes caused by sensor noise, adverse weather, and dynamic lighting can degrade perception, leading to suboptimal control decisions. Existing approaches, including domain adaptation and adversarial training, improve robustness but struggle to generalize to unseen corruptions while introducing computational overhead. To address this challenge, we propose a real-time image repair module that restores corrupted images before they are used by the controller. Our method leverages generative adversarial models, specifically CycleGAN and pix2pix, for image repair. CycleGAN enables unpaired image-to-image translation to adapt to novel corruptions, while pix2pix exploits paired image data when available to improve the quality. To ensure alignment with control performance, we introduce a control-focused loss function that prioritizes perceptual consistency in repaired images. We evaluated our method in a simulated autonomous racing environment with various visual corruptions. The results show that our approach significantly improves performance compared to baselines, mitigating distribution shift and enhancing controller reliability.
@@
Demonstration video: youtu.be/C3-WlZpYBm8
@@ I INTRODUCTION
Vision-based control has achieved significant progress, particularly in racing applications [1]. Despite these advancements, real-world deployment remains challenging due to image distribution shifts [2] caused by sensor noise, adverse weather conditions (such as rain, fog, and snow), rapid lighting changes, and adversarial perturbations. Such factors can lead to severe image corruption, compromising the stability and reliability of visual control systems. In practice, most existing vision-based methods assume that the training data accurately reflects the test environment. However, this assumption often fails, resulting in subpar decision-making and degraded performance.
@@
This paper addresses the problem of image repair for robust control. We introduce a real-time image repair module that seamlessly integrates into vision-based control systems. This module repairs degraded images at test time without any prior knowledge of them, thereby reducing the distribution gap between the training and testing environments. Our method leverages recent advances in GANs, specifically CycleGAN [9] and pix2pix [10], to perform real-time image restoration. CycleGAN enables unpaired image-to-image translation, allowing adaptation to unseen corruptions without requiring paired training data, while pix2pix utilizes paired data to learn a direct mapping between corrupted and clean images, leading to more precise restorations when such data is available.
By utilizing these image translation techniques, our approach balances computational efficiency and repair quality, ensuring minimal latency while preserving essential perception features for reliable control. To train our module, we introduce a control-focused loss function that ensures the repaired images lead to accurate control outputs, directly improving performance under adverse conditions.
@@
  • A real-time image repair module that integrates into vision-based control systems, enhancing robustness against image corruption.
@@ II PROBLEM STATEMENT
  • Vision-based control module $h:\mathbf{Y}\rightarrow\mathbf{U}$, which takes an image as input and is realized by a neural network
@@
Given a training dataset $\mathcal{D}_{train}$ of paired images and control actions, where the images are drawn i.i.d from the image distribution $P_{train}$ and the actions follow the center line of the track, we define an image-based neural network controller $h$, which translates image observations $y$ into control actions $u$. During operation, we have a testing image distribution $P_{test}$, where images may undergo corruptions and disturbances. To formalize the relationship between the testing and training images, we assume that test images in $P_{test}$ are generated from clean images in $P_{train}$ via an unknown random noise $\eta$ from an unknown distribution $P_{\eta}$. In our experiments, the corruptions we consider include darkness, salt-and-pepper noise, rain, fog, and snow, which we denote as $\eta\in\{\eta_{\text{dark}},\eta_{\text{salt}},\eta_{\text{rain}},\eta_{\text{fog}},\eta_{\text{snow}}\}$. Examples of these corruptions are shown in Figure 1. Then we consider two settings of image generation:
@@
We define an image repair model, $r:\mathbf{Y}\rightarrow\mathbf{Y}$, that translates images from the testing domain to the training domain without explicit knowledge of the corruption functions $c$ and $C$ or noises $\eta$ or $P_{\eta}$. Intuitively, the repair model $r$ aims to shift a corrupted image $\hat{y}$ closer to training distribution $P_{train}$ before passing it to the controller:
$y^{\prime}=r(\hat{y}),$ (4)
where $y^{\prime}$ is the repaired image. Our architecture is shown in Figure 2. Images from the observation model, $o$, are passed to our repair model $r$, to obtain repaired images. These repaired images are then used by the controller $h$ to predict control actions.
@@
For the autonomous car racing setting, we calculate this difference in trajectory by comparing mean-squared cross-track error (CTE) values with respect to the center line of the track.
@@ III IMAGE REPAIR FOR CONTROL
This section introduces the repair techniques that we experimented with and then describes our control-focused training for generative models.
@@ III-A Repair Techniques
Adversarial restoration: Much like in adversarial training, many restoration methods opt for learning-based approaches. For instance, PixelDefend [15] and generative adversarial denoising [16] reconstruct perturbed images to restore clean data representations, randomized smoothing [17] certifies robustness by averaging predictions over noisy variations of an input, and feature denoising networks [18] integrate adversarial noise suppression within convolutional architectures. However, these methods also come with increased computational complexity, and their efficacy often depends on the strength and nature of adversarial attacks.
@@
where $r_{\text{LR}}$ is the Lucy Richardson Deconvolution approach, $r_{\text{VB}}$ is the Variational Bayesian approach, $\hat{y}$ is the original image from $P_{test}$, and $y^{\prime}$ is the repaired image that is then passed to the controller as follows:
@@
We implement three generative approaches — VAEs, CycleGAN, and pix2pix — and compare their performance in Section IV. We prepend our generative repair model $r$ to $h$, repairing corrupted images $\hat{y}\in P_{test}$ and passing them to $h$:
@@ III-B Generative Model Training
For pix2pix, the training parallels that of CycleGAN, but with the requirement of having a paired dataset of normal and corrupted images. We train for 30 epochs and then take the generator as our repair model. Much like for CycleGAN, we implement an additional controller loss $\mathcal{L}_{GS}$, representing the L1 loss between the predicted control action on normal images $y$ and predicted images $y^{\prime}$ from the generator. The loss is then added to the existing pix2pix generator loss:
@@ IV EXPERIMENTAL EVALUATION
Config Approach Paired-ness Controller Type Normal Brightness Salt/Pepper Rain Fog Snow Unseen All
@@
- Lucy Richardson - - 2.957 3.148 2.459 2.088 1.969 1.923 2.362 2.471
- Variational Bayes - - 2.158 3.078 2.552 2.733 2.523 2.275 2.646 2.571
@@
VAE - - 1.395 3.147 2.088 1.838 1.942 2.220 2.085 2.171
CycleGAN Unpaired tr: $\varnothing$, te: $h$ 1.611 1.499 2.559 1.424 1.178 1.965 1.620 1.764
CycleGAN Unpaired tr: $\varnothing$, te: $h_{adv}$ 1.282 1.525 1.726 1.533 1.475 2.306 1.936 1.673
CycleGAN Paired tr: $h$, te: $h$ 1.242 1.353 1.345 1.392 1.087 1.833 1.507 1.394
CycleGAN Paired tr: $h$, te: $h_{adv}$ 1.519 1.392 2.234 1.396 2.347 2.364 2.355 1.927
CycleGAN Paired tr: $h_{adv}$, te: $h$ 1.466 1.709 1.885 1.405 2.069 1.956 2.013 1.765
CycleGAN Paired tr: $h_{adv}$, te: $h_{adv}$ 1.502 1.763 1.602 1.663 1.480 2.320 1.946 1.745
pix2pix Paired tr: $\varnothing$, te: $h$ 1.648 1.403 1.552 1.451 1.415 2.003 1.734 1.605
pix2pix Paired tr: $\varnothing$, te: $h_{adv}$ 1.588 1.767 1.628 2.323 1.464 2.154 1.842 1.847
pix2pix Paired tr: $h$, te: $h$ 1.984 1.950 1.966 1.922 1.602 2.028 1.936 1.951
pix2pix Paired tr: $h$, te: $h_{adv}$ 1.635 1.552 2.261 2.312 2.361 2.276 2.319 2.095
pix2pix Paired tr: $h_{adv}$, te: $h$ 1.817 1.846 1.931 1.997 2.121 1.959 2.042 1.948
pix2pix Paired tr: $h_{adv}$, te: $h_{adv}$ 1.655 2.189 1.534 2.149 2.254 2.124 2.190 2.004
@@
VAE - - 1.393 3.147 2.146 1.923 1.935 2.071 1.998 2.167
CycleGAN Unpaired tr: $\varnothing$, te: $h$ 1.500 1.446 1.723 1.264 1.266 2.189 1.788 1.597
CycleGAN Unpaired tr: $\varnothing$, te: $h_{adv}$ 1.405 1.419 1.516 1.639 2.239 2.106 1.887 1.752
CycleGAN Paired tr: $h$, te: $h$ 1.570 1.506 1.117 1.513 1.338 1.995 1.770 1.530
CycleGAN Paired tr: $h$, te: $h_{adv}$ 1.507 1.567 1.412 1.668 1.639 2.075 1.883 1.658
CycleGAN Paired tr: $h_{adv}$, te: $h$ 1.265 2.332 2.782 2.580 2.830 1.869 2.253 2.343
CycleGAN Paired tr: $h_{adv}$, te: $h_{adv}$ 1.455 2.628 2.265 2.297 2.702 2.166 2.233 2.289
pix2pix Paired tr: $\varnothing$, te: $h$ 2.652 2.250 1.853 1.586 1.673 1.865 1.731 2.013
pix2pix Paired tr: $\varnothing$, te: $h_{adv}$ 1.594 1.719 1.676 1.888 2.124 2.071 1.981 1.856
pix2pix Paired tr: $h$, te: $h$ 1.840 1.548 1.540 1.529 1.454 1.919 1.784 1.656
pix2pix Paired tr: $h$, te: $h_{adv}$ 2.063 1.614 1.643 1.809 1.761 2.085 1.952 1.838
pix2pix Paired tr: $h_{adv}$, te: $h$ 1.796 1.804 1.682 1.574 1.406 1.341 1.523 1.643
pix2pix Paired tr: $h_{adv}$, te: $h_{adv}$ 1.801 1.875 1.743 1.817 1.621 2.066 1.946 1.826
TABLE II: Average Inference Time per Image
Model Time (ms)
Original Controller ($h$) 1.76 $\pm$ 0.84
Lucy Richardson 0.96 $\pm$ 0.20
Variational Bayes 5.38 $\pm$ 1.49
VAE 2.52 $\pm$ 3.57
CycleGAN 23.86 $\pm$ 6.40
pix2pix 6.88 $\pm$ 2.47
Experimental Setup. The primary objective is to assess how well different repair techniques restore corrupted images and improve control performance under various disturbances. To evaluate the effectiveness of our image repair methods in vision-based autonomous racing, we conduct experiments in a simulated environment for DonkeyCar racing (see https://www.donkeycar.com/). Simulation is chosen over real-world deployment due to its controlled setting, reproducibility, ability to systematically introduce diverse corruptions, and an easy-to-quantify autonomous racing task. On the other hand, open datasets like nuScenes and Waymo Open do not allow counterfactual executions under visual corruptions, whereas implementing repeatable paired corruptions in a physical setting can be very effortful. The implementation of our evaluation can be found in an online repository: https://github.com/Trustworthy-Engineered-Autonomy-Lab/generalizable-image-repair.
@@
We propose using CycleGAN and pix2pix. For CycleGAN, we experiment with both unpaired and paired training with a controller loss. The unpaired CycleGAN learns a mapping between $P_{test}$ (corrupted images) and $P_{train}$ (clean images) without explicit supervision. In contrast, the paired CycleGAN is trained with corrupted-clean image pairs and incorporates an additional loss function to align repaired images with optimal control actions. Finally, we train a pix2pix model on paired corrupted-clean images, both with and without a controller loss, to examine its effectiveness compared to CycleGAN-based methods.
@@
Overall, these results confirm the effectiveness of integrating generative repair models for mitigating visual corruptions. Notably, incorporating a controller loss during CycleGAN training significantly enhances repair quality, leading to more stable and reliable control under challenging conditions. This highlights the importance of aligning image repair objectives with downstream control performance, rather than relying solely on visual reconstruction quality.
Inference Time. To evaluate the run-time feasibility, we report the average inference time and standard deviation in Table II. Each model was evaluated on 1000 images from the test set using an NVIDIA L4 GPU. Overall, Lucy Richardson and the original controller exhibit the fastest inference, followed by VAE and Variational Bayes. In contrast, pix2pix and CycleGAN are notably slower yet still feasible, indicating a trade-off between performance and computing cost.
@@ V-C Adversarial Robustness
For robustness against adversarial perturbations, adversarial training (AT) [36] uses adversarial examples in training. However, as previously discussed, AT still struggles to generalize to adversarial examples outside of the training set. While the controller performs well on the three known corruptions, it struggles to generalize to the other two unseen noises.
Instead of modifying training, adversarial purification removes perturbations from input images. To compare our method to purification approaches, we tested Lucy Richardson Deconvolution and Variational Bayes, which proved detrimental to controller performance on both normal and corrupted images. Thus, classical computer vision approaches struggle to handle complex and unstructured noises, such as those seen in the real world (rain, fog, snow).